# Kinetic localization via Poincaré-type inequalities and applications to the condensation of Bose gases

### v1

@@ 1. Introduction
for all $u\in H^{1}(\Omega)$ (see, e.g., [24, Theorem 8.11]). Here, $\left\langle f\right\rangle_{A}:=|A|^{-1}\int_{A}f$ is the integral mean of $f$, with $|A|$ the Lebesgue measure of $A\subset\mathbb{R}^{d}$. For convenience, we write $\int_{A}f$ instead of $\int_{A}f\,\mathrm{d}x$. Equivalently, (1) can be interpreted as the existence of a spectral gap of size $C_{\Omega}^{-1}$ between the first eigenvalue and the second eigenvalue of the Neumann Laplacian $\Delta^{\mathrm{Neu}}_{\Omega}$ on $L^{2}(\Omega)$, namely
@@
Let us mention that the need for a Poincaré inequality in the context of Bose–Einstein condensation is well known; see, for instance, the fundamental work [28]. However, the Poincaré inequality derived in our paper is simpler than the existing ones, while still being useful for understanding the Bose gas. We hope that our simplified approach will stimulate further studies of the localization argument in the context of Bose–Einstein condensation in the future.
@@ Acknowledgments
We would like to thank Zhenfu Wang for many helpful discussions. J.J. Chong thanks Søren Fournais for discussions on the localization methods used in his work, their connection to the IMS inequality, and their history. H. Liang thanks Zhibin Gong for insightful discussions. This work was partially supported by the National Key R&D Program of China (Project No. 2024YFA1015500, J.J. Chong and H. Liang) and by the European Research Council through the ERC Consolidator Grant RAMBAS (Project No. 10104424, P.T. Nam).
@@ 2. Main Results
2.1. Generalized Poincare inequality for cubes
@@ 2.2. Bose–Einstein condensation of dilute Bose gases
We say that a state $\Psi\in L_{s}^{2}(\Lambda^{N})$ exhibits 100% Bose–Einstein condensation (complete BEC) on the constant function $u_{0}=1\in L^{2}(\Lambda)$ if
@@
Theorem 2 (Complete BEC).
@@
The LHY formula suggests that, for the ground state, the “optimal” second-order term may scale as $N^{5\kappa/2}$. Here, however, we instead assume the condition $N^{5\kappa/2+(2-3\kappa)/4}$, which equals $\sqrt{N}\gg 1$ when $\kappa=0$. This specific constraint, $N^{5\kappa/2+(2-3\kappa)/4}$, arises from the error estimate in Proposition A.1 with $S(\cdot)=\sqrt{\cdot}$. By treating the “completing-the-square” step more carefully (and taking into account the cubic terms), one may expect to improve Proposition A.1 and choose $S(\cdot)=\log(\cdot)$, as already done in [5], using a more involved method based on unitary transformations. In that case, complete BEC can be obtained for $\kappa\in(0,\frac{2}{7}-\varepsilon)$, provided that the stricter upper bound $N^{5\kappa/2+\varepsilon}$ holds for the second-order term, with $\varepsilon>0$ small arbitarily. That is, applying the result of [5, Theorem 1.1] in our proof of Theorem 2 yields the following result.
@@
with some constant $C_{0}>0$ independent of $N$. Then we obtain complete BEC (14).
@@ 2.3. Remarks on previous results
In the case of beyond-GP scaling, the proof of BEC—although not explicitly stated—was already implied by the work [25] for $\kappa\in[0,\frac{1}{10})$ (see [31, Chapter 7]). A different approach, based on unitary renormalizations, was developed in [1] to prove BEC for approximate ground states in the regime $\kappa\in[0,\frac{1}{43})$; this was later revisited and improved in [7] to cover the regime $\kappa\in[0,\frac{1}{20})$. Currently, the best available result in terms of the range of $\kappa$ is given in [15] for the regime $\kappa\in[0,\frac{2}{5})$.
Localization methods play a central role in all existing proofs of BEC. In particular, the best-known result in [15] was obtained using a subtle localization argument that had been developed earlier to derive the Lee–Huang–Yang formula [10, 9, 18, 19] (building on even earlier works [13, 29, 30]). Historically, a more natural localization method based on Neumann box localization was used in [32], which inspired the proof of BEC in [25] and the generalization of the Poincare inequality in [28]. More recently, new techniques concerning Neumann localization have been developed in [20], leading to an extension of the Lee–Huang–Yang formula to positive temperatures.
In the present paper, we combine the techniques from [35, 20] with Inequality (5) in Theorem 1 to provide a simplified, self-contained proof of BEC for $\kappa\in(0,\frac{2}{11})$. More precisely, in Section A, we establish in Proposition A.1 the simplest case of Theorem 2, namely when $\kappa=0$ and the scattering length of $V$ is sufficiently small, by combining techniques from [35, 20]. Then, in Section 3, we prove Theorem 1, which is of independent interest. Finally, in Section 4, we explain how Theorem 1 can be used to implement a localization argument, allowing us to deduce Theorem 2 from Proposition A.1. Moreover, Theorem 3 represents a minor modification of the proof of Theorem 2 but relies heavily on [5, Theorem 1.1].
3. Poincare-type Inequality
@@
A proof of (21), with slight modification, can be founded in [33, Proposition 3] (see also [12, Theorem 2.8]), where we have used the fact that the Cheeger constant of $\llbracket M\rrbracket^{d}$ is inversely proportional to $M$. Combining (20), (21), and the fact $M=1/\ell$ yields
@@ 4.2. Proof of Theorem 2
We follow a similar approach as in the proof given in [15, Theorem 2.1]. Since all these operators are unitary equivalent, it suffices to consider $i=1$. Notice that $\mathcal{H}_{\rho_{\mu}}(\Lambda_{1})$ commutes with the particle number operator, so it suffices to establish the lower bound for $\left\langle\Psi,(\mathcal{H}_{\rho_{\mu}}(\Lambda_{1}))_{N}\Psi\right\rangle$ for $\Psi\in L^{2}_{s}(\Lambda_{1}^{N})$ for some arbitrary particle number $N$. Since $\rho_{\mu}\ell^{3}=K^{-3}(\rho_{\mu}\mathfrak{a}^{3})^{-1/2}\gg 1$, we partition the $N$ particles into groups with $\rho_{\mu}\ell^{3}$ order number of particles. Let
@@
which can be obtained via a standard localization method (see e.g. [20, Eq. (9.6)]). Hence, it suffices to reduce the analysis to $(\mathcal{H}_{\rho_{\mu}}(\Lambda_{1}))_{n}$ with $n\leq 4\rho_{\mu}\ell^{3}$, i.e. $\frac{n\mathfrak{a}}{\ell}\leq\frac{1}{100}.$
@@ A.1. Modified scattering function and Neumann symmetrization
The proof can be found in [20, Lemma 3.2]. Here, we define the Fourier transform $f$ by $\hat{f}(p)=\int_{\mathbb{R}^{3}}f(x)e^{-ix\cdot p}\,\mathrm{d}x.$
@@
The proof is exactly the same as in [20, Lemma 3.3], with $n+K(x,y)=nF(x,y)$ in our notations. ∎
@@
Recall the simple case of Bogoliubov’s method [29, Theorem 6.3], we have

### v2

@@ 1. Introduction
for all $u\in H^{1}(\Omega)$ (see, e.g., [26, Theorem 8.11]). Here, $\left\langle f\right\rangle_{A}:=|A|^{-1}\int_{A}f$ is the integral mean of $f$, with $|A|$ the Lebesgue measure of $A\subset\mathbb{R}^{d}$. For convenience, we write $\int_{A}f$ instead of $\int_{A}f\,\mathrm{d}x$. Equivalently, (1) can be interpreted as the existence of a spectral gap of size $C_{\Omega}^{-1}$ between the first eigenvalue and the second eigenvalue of the Neumann Laplacian $\Delta^{\mathrm{Neu}}_{\Omega}$ on $L^{2}(\Omega)$, namely
@@
The usefulness of a Poincaré inequality in the context of BEC is well known; see, for instance, the fundamental work [30]. However, the Poincaré inequality derived in our paper is simpler than existing ones, while remaining effective for analyzing the Bose gas. We hope this simplified approach will stimulate further research on localization arguments in BEC.
@@ Acknowledgments
We would like to thank Zhenfu Wang for many helpful discussions. J.J. Chong thanks Søren Fournais for discussions on the localization methods used in his work, their connection to the IMS inequality, and their history. H. Liang thanks Zhibin Gong for insightful discussions. P. T. Nam thanks Christian Brennecke for insightful remarks comparing various proofs of Bose–Einstein condensation. This work was partially supported by the National Key R&D Program of China (Project No. 2024YFA1015500, J.J. Chong and H. Liang) and by the European Research Council through the ERC Consolidator Grant RAMBAS (Project No. 10104424, P.T. Nam).
@@ 2. Main Results
2.1. Generalized Poincaré inequality for cubes
@@
Remark 3.
We note that the Poincaré inequality (4) is closely related, albeit in completely different settings and form, to the multiscale Poincaré inequality, which is used in quantitative homogenization theory [2]. Similar estimates have also been developed to establish quantitative homogenization for interacting particle systems [21].
@@ 2.2. Bose–Einstein condensation of dilute Bose gases
We say that a state $\Psi\in L_{s}^{2}(\Lambda^{N})$ exhibits Bose–Einstein condensation (BEC) on the constant function $u_{0}=1\in L^{2}(\Lambda)$ if
@@
Theorem 2 (Bose–Einstein condensation).
@@
The LHY formula suggests that, for the ground state, the “optimal” second-order term may scale as $N^{5\kappa/2}$. Here, however, we instead assume the condition $N^{5\kappa/2+(2-3\kappa)/4}$, which equals $\sqrt{N}\gg 1$ when $\kappa=0$. This specific constraint, $N^{5\kappa/2+(2-3\kappa)/4}$, arises from the error estimate in Proposition A.1 with $S(\cdot)=\sqrt{\cdot}$. By treating the “completing-the-square” step more carefully (and taking into account the cubic terms), one may expect to improve Proposition A.1 and choose $S(\cdot)=\log(\cdot)$, as already done in [6], using a more involved method based on unitary transformations. In that case, complete BEC can be obtained for $\kappa\in(0,\frac{2}{7}-\varepsilon)$, provided that the stricter upper bound $N^{5\kappa/2+\varepsilon}$ holds for the second-order term, with $\varepsilon>0$ small arbitarily. That is, applying the result of [6, Theorem 1.1] in our proof of Theorem 2 yields the following result.
@@
with some constant $C_{0}>0$ independent of $N$. Then we obtain BEC (14).
@@ 2.3. Remarks on previous results
In the case of beyond-GP scaling, the proof of BEC—although not explicitly stated—was already implied by the work [27] for $\kappa\in[0,\frac{1}{10})$ (see [33, Chapter 7]). A different approach, based on unitary renormalizations, was developed in [1] to prove BEC for approximate ground states in the regime $\kappa\in[0,\frac{1}{43})$; this was later revisited and improved in [8] to cover the regime $\kappa\in[0,\frac{1}{20})$. Currently, the best available result in terms of the range of $\kappa$ is given in [16] for the regime $\kappa\in[0,\frac{2}{5})$.
Localization methods are typically helpful in proofs of Bose–Einstein condensation (BEC), although some works do not rely on them; see, for instance, [1, 8] for alternative approaches that require neither $x$-localization nor the smallness of the interaction. (On the technical side, the approaches in [3, 37, 6] also avoid localization in the actual proof of BEC, but these works rely on a smallness assumption on the interaction, which is incompatible with going beyond the Gross–Pitaevskii regime.) The best-known result in [16], in particular, was obtained using a subtle localization argument developed to derive the Lee–Huang–Yang formula [11, 10, 19, 20], building on even earlier works [14, 31, 32]. Historically, a more natural localization method based on Neumann box confinement was used in [34], which inspired the proof of BEC in [27] and the generalization of the Poincaré inequality in [30]. More recently, new techniques concerning Neumann localization have been developed in [22], leading to an extension of the Lee–Huang–Yang formula to positive temperatures.
In the present paper, we combine the techniques from [37, 22] with Inequality (5) in Theorem 1 to provide a simplified, self-contained proof of BEC for $\kappa\in(0,\frac{2}{11})$. More precisely, in Section 3, we prove the Poincaré-type inequality in Theorem 1, which is of independent interest. Then, in Section 4, we explain how Theorem 1 can be used to implement a localization argument, allowing us to deduce BEC beyond the GP regime from a quantitative lower bound of the Hamiltonian imposed with Neumann boundary condition within the GP regime, namely when $\kappa=0$. To make our proof self-contained, we establish in Proposition A.1 in Appendix A a quick proof of such lower bound when the scattering length of $V$ is sufficiently small, by combining techniques from [37, 22]. This, together with the localization argument, implies BEC for $\kappa\in(0,\frac{2}{11})$. To improve this result (see Theorem 3), a refinement of Proposition A.1 is needed. Theorem 3 represents a minor modification of the proof of Theorem 2 but relies heavily on [6, Theorem 1.1].
3. Poincaré-type Inequality
@@
A proof of (21), with slight modification, can be founded in [35, Proposition 3] (see also [13, Theorem 2.8]), where we have used the fact that the Cheeger constant of $\llbracket M\rrbracket^{d}$ is inversely proportional to $M$. Combining (20), (21), and the fact $M=1/\ell$ yields
@@ 4.2. Proof of Theorem 2
We follow a similar approach as in the proof given in [16, Theorem 2.1]. Since all these operators are unitary equivalent, it suffices to consider $i=1$. Notice that $\mathcal{H}_{\rho_{\mu}}(\Lambda_{1})$ commutes with the particle number operator, so it suffices to establish the lower bound for $\left\langle\Psi,(\mathcal{H}_{\rho_{\mu}}(\Lambda_{1}))_{N}\Psi\right\rangle$ for $\Psi\in L^{2}_{s}(\Lambda_{1}^{N})$ for some arbitrary particle number $N$. Since $\rho_{\mu}\ell^{3}=K^{-3}(\rho_{\mu}\mathfrak{a}^{3})^{-1/2}\gg 1$, we partition the $N$ particles into groups with $\rho_{\mu}\ell^{3}$ order number of particles. Let
@@
which can be obtained via a standard localization method (see e.g. [22, Eq. (9.6)]). Hence, it suffices to reduce the analysis to $(\mathcal{H}_{\rho_{\mu}}(\Lambda_{1}))_{n}$ with $n\leq 4\rho_{\mu}\ell^{3}$, i.e. $\frac{n\mathfrak{a}}{\ell}\leq\frac{1}{100}.$
@@ A.1. Modified scattering function and Neumann symmetrization
The proof can be found in [22, Lemma 3.2]. Here, we define the Fourier transform $f$ by $\hat{f}(p)=\int_{\mathbb{R}^{3}}f(x)e^{-ix\cdot p}\,\mathrm{d}x.$
@@
The proof is exactly the same as in [22, Lemma 3.3], with $n+K(x,y)=nF(x,y)$ in our notations. ∎
@@
Recall the simple case of Bogoliubov’s method [31, Theorem 6.3], we have